### v1

Abstract
In scientific research and its application, scientific literature analysis is crucial as it allows researchers to build on the work of others. However, the fast growth of scientific knowledge has led to a massive increase in scholarly articles, making in-depth literature analysis increasingly challenging and time-consuming. The emergence of Large Language Models (LLMs) has offered a new way to address this challenge. Known for their strong abilities in summarizing texts, LLMs are seen as a potential tool to improve the analysis of scientific literature. However, existing LLMs have their own limits. Scientific literature often includes a wide range of multimodal elements, such as molecular structure, tables, and charts, which are hard for text-focused LLMs to understand and analyze. This issue points to the urgent need for new solutions that can fully understand and analyze multimodal content in scientific literature. To answer this demand, we present Uni-SMART (Universal Science Multimodal Analysis and Research Transformer), an innovative model designed for in-depth understanding of multimodal scientific literature. Through rigorous quantitative evaluation across several domains, Uni-SMART demonstrates superior performance over leading text-focused LLMs. Furthermore, our exploration extends to practical applications, including patent infringement detection and nuanced analysis of charts. These applications not only highlight Uni-SMART’s adaptability but also its potential to revolutionize how we interact with scientific literature.
1 Introduction
Scientific literature, encompassing patents and academic papers, constitutes a rich science data resource, including but not limited to drug properties and activities, reaction pathways, and manufacturing processes. However, extracting target information from this extensive corpus is a laborious and time-intensive task. It necessitates meticulous manual review, analysis, and extraction – processes that are inherently slow and prone to human error [1, 2]. To enhance the efficiency of information retrieval, specialized databases like Sci-Finder [3] and Reaxys [4] have been developed. However, their utility is constrained to document retrieval for molecule and reaction queries, lacking the capabilities of information extraction and knowledge comprehension to function as domain assistants. Consequently, users must still engage in the tedious tasks of reading and analyzing the retrieved documents to extract definitive answers. This limitation poses a significant bottleneck in the utilization of scientific data, hindering research progress and the rapid application of discoveries. Thus, researchers and practitioners require an intelligent navigator that can swiftly guide through the complexities of the latest scientific data, identify relevant information with precision, and present it in a digestible format.
Figure 1: Uni-SMART overview. It interprets multimodal data (e.g., texts, molecule structures, chemical reactions, charts, and tables), thereby facilitating a broad range of applications such as patent infringement analysis, chart analysis, and more.
The emergence of Large Language Models (LLMs), represented by LLaMA [5], Gemini [6], and GPT [7, 8, 9], has marked a significant milestone in the evolution of natural language processing. These models have revolutionized the extraction of textual information from documents, enabling direct responses to queries using the extracted content. Despite their proficiency, current LLMs are primarily designed for text extraction and often struggle with the multimodal aspects inherent in scientific literature, which include a large number of tables, charts, and reactions. The extraction and interpretation of such multimodal data require an understanding that beyond texts and delves into visual and structural contents.
To address these challenges, we developed Uni-SMART (Universal Science Multimodal Analysis and Research Transformer), which extends the capabilities of LLMs beyond text, allowing for the interpretation of the multimodal content that is crucial in scientific literature. As illustrated in Figure 1, It is designed to recognize and analyze multimodal data, such as molecule structures, chemical reactions, charts, and tables, alongside textual content, facilitating a comprehensive understanding of scientific literature. Such ability not only augments automated and precise information extraction but also enriches the interaction between researchers and the vast expanse of scientific knowledge.
To rigorously assess the multimodal capabilities of Uni-SMART, a comparative analysis was conducted against leading LLMs, such as Gemini, GPT-3.5, and GPT-4. Our assessment targets several data types critical to the comprehension of scientific documents: tables, charts, molecular structures, and chemical reactions. The results demonstrate Uni-SMART’s superior performance in all tested areas, especially in understanding and analyzing complex multimodal contents, thus highlighting its potential as a helpful assistant for scientific literature analysis.
In the following sections, Section 2 details Uni-SMART’s data sources and iterative training approach. Subsequently, Section 3 – Evaluation, presents detailed comparisons of Uni-SMART with leading LLMs across a variety of modalities, showcasing its advanced capabilities in multimodal data interpretation. Then, Section 4 – Application showcases some specific applications of Uni-SMART in the scientific domain. Finally, Section 5 discusses Uni-SMART’s limitations and future research avenues, along with its potential impact on scientific research and technological advancement.
2 Method
As depicted in Figure 2, Uni-SMART sources training data from a wide range of scientific literature from global patents, news articles, scientific publications, and market reports. In particular, It adopts a cyclical, iterative approach to enhance its multimodal understanding capabilities, comprising the following key components:
  • Multimodal Learning: During the initialization phase, the model is trained with a limited set of multimodal data to recognize and extract diverse information elements from scientific literature, outputting these in a sequential format that incorporates both textual and multimodal data.
  • LLM SFT: The output sequences, along with corresponding question-answer pairs, are utilized for supervised fine-tuning (SFT) of LLM, boosting its proficiency in handling multimodal content.
  • User Feedback: The SFT-enhanced LLM is deployed in real applications, facilitating the collection of user feedback. Samples receiving positive feedback are subsequently filtered and incorporated into the data enhancement, while those with negative feedback are subject to expert annotation before being integrated into the data enhancement process.
  • Expert Annotation: Samples with negative feedback are meticulously annotated by human experts. This step ensures that the model learns from its mistakes, with semi-automated tools assisting in this process to enhance efficiency. Negative feedback is examined to identify challenges in multimodal recognition or the language model’s reasoning abilities. This analysis helps determine which multimodal elements and types of question are more challenging for the model, facilitating focused improvements.
  • Data Enhancement: Finally, the annotated data, along with partial samples that received positive feedback, are added to the training dataset for data enhancement. This iterative process progressively improves Uni-SMART’s performance, enriching its training dataset and enhancing its multimodal understanding capabilities.
Such cyclical pipeline significantly enhances Uni-SMART’s performance in a variety of challenging tasks, such as information extraction, complex element identification, scientific literature understanding/analysis, and multimodal understanding/reasoning.
Figure 2: Uni-SMART model architecture.
3 Evaluation
In this section, we perform a detailed quantitative evaluation on the capabilities of Uni-SMART and leading LLMs across various modalities. Table 1 presents the statistics of the benchmarks. For details of the specific evaluation tasks, please refer to SciAssess [10]. And the experimental setup for each model is detailed in Appendix A.
Table 1: Dataset statistics.
Modality
Domain
Task
Question #
Table
Alloy Materials
Composition Extraction
55
Drug Discovery
Affinity Data Extraction
15
Organic Materials
Electrolyte Solubility data extraction
8
Electrolyte Table QA
48
Polymer Property Extraction
15
OLED Property Extraction
13
Chart
Alloy Materials
Alloy ChartQA
15
Biology
Biology ChartQA
15
Drug Discovery
Drug ChartQA
15
Organic Materials
Polymer ChartQA
15
Molecule
Drug Discovery
Affinity Data Extraction
15
Tag to Molecule
41
Markush to Molecule
9
Organic Materials
OLED Property Extraction
13
Chemical Reaction
Drug Discovery
Reaction QA
15
Organic Materials
Reaction Mechanism QA
22
3.1 Table
Tables play a pivotal role in scientific literature, presenting complex data and findings in a highly structured manner and thereby contributing significantly to scientific discovery and the dissemination of knowledge [11]. They facilitate the intuitive display of experimental data and enable the efficient summarization and comparison of research outcomes, becoming an indispensable component of scientific investigation. Consequently, enhancing the capability to understand tables is crucial for the automated processing and analysis of scientific documents. The utilization of table data spans a wide array of applications, such as trend analysis, which can reveal developmental trajectories in research fields, and comparative studies, which can elucidate differences in experimental outcomes under varying research methodologies or conditions [12].
To assess the table understanding capabilities of our model, Uni-SMART, compared to other leading models, we designed a diverse set of tasks across different domains. These tasks were specifically tailored to evaluate how well each model could interpret and extract information from tables, with the results presented in Table 2.
In our thorough assessment, the Uni-SMART showcased its exceptional proficiency in understanding and extracting table data from the scientific literature. Among six assessed tasks, it surpassed other models in four and delivered a competitive performance in the remainder. Notably, in the "Electrolyte Table QA" and "Polymer Property Extraction" tasks, Uni-SMART achieved Value Recall scores of 0.674 and 0.869, respectively, significantly outperforming its counterparts. These results highlight Uni-SMART’s exceptional skill in processing table data. However, Uni-SMART’s performance in the "Composition Extraction" task for Alloy Materials showed no significant difference compared to other models, only slightly leading. To further enhance its effectiveness, a possible approach could be to strengthen the model’s ability to learn domain-specific knowledge. Moreover, all models under-performed in the "Affinity Data Extraction" task within Drug Discovery and the "Electrolyte Solubility data extraction" task in Organic Materials, indicating the high complexity of these tasks. Such challenges might be attributed to the diversity of table formats and the presence of complex nested structures, such as tables embedding molecular structures. To navigate these difficulties, future directions could involve the development of more sophisticated table data comprehension technologies and the refinement of data preprocessing methods.
Table 2: Evaluation on table data.
Domain
Task
Metric
Uni-SMART
GPT-4
GPT-3.5
Gemini
Alloy Materials
Composition Extraction
Value Recall
0.451
0.420
0.437
0.443
Drug Discovery
Affinity Data Extraction
Value Recall
0.320
0.274
0.359
0.175
Organic
Materials
Electrolyte Solubility data extraction
Value Recall
0.282
0.365
0.393
0.390
Electrolyte Table  QA
Value Recall
0.674
0.163
0.140
0.233
Polymer Property Extraction
Value Recall
0.869
0.762
0.286
0.506
OLED Property Extraction
Value Recall
0.526
0.520
0.479
0.394
3.2 Chart
Charts are essential tools in scientific literature, offering a visual representation of data that can significantly enhance the comprehension and communication of complex information. By succinctly illustrating trends, comparisons, and patterns, charts enable researchers to convey their findings more effectively and intuitively [13]. Therefore, the ability to accurately interpret and analyze charts is vital for the automated processing and understanding of scientific documents.
To assess the chart interpretation capabilities of Uni-SMART, we conducted a series of ChartQA tasks spanning various scientific domains. These tasks rigorously evaluated models’ ability to analyze and clarify data presented in charts, with a particular focus on identifying trends and extracting meaningful insights.
Analysis of the results presented in Table 3 indicates that Uni-SMART significantly outperforms existing models like GPT-4, GPT-3.5, and Gemini in ChartQA tasks across diverse scientific domains, particularly excelling in the areas of alloy materials and organic materials, with leading scores of 0.667 and 0.733, respectively. However, despite Uni-SMART’s leading position, its advantage narrows in the domains of drug discovery and biology, where all models struggle to achieve high performance. This limitation is likely due to the complexity and specificity of charts in these fields, which demand higher levels of understanding and adaptability from the models. To overcome these challenges, future work should focus on augmenting the training dataset with a wider and more sophisticated assortment of chart data, aiming to enhance the model’s analytical depth and accuracy. Furthermore, exploring refined model architectures and specialized training methodologies, specifically designed to meet the complex requirements of chart analysis, is crucial for significantly elevating Uni-SMART’s performance in such demanding tasks.
Table 3: Evaluation on chart data.
Domain
Task
Metric
Uni-SMART
GPT-4
GPT-3.5
Gemini
Alloy Materials
Alloy ChartQA
Accuracy
0.667
0.400
0.267
0.333
Biology
Biology ChartQA
Accuracy
0.467
0.333
0.400
0.400
Drug Discovery
Drug ChartQA
Accuracy
0.467
0.400
0.333
0.400
Organic Materials
Polymer ChartQA
Accuracy
0.733
0.667
0.333
0.007
3.3 Molecular Structure
In scientific literature, molecular structures hold fundamental importance, especially within fields such as chemistry, pharmacology, and biology. The ability to comprehend molecular structures is crucial for analyzing research outcomes, predicting chemical behaviours, and innovating novel compounds [14, 15]. Deciphering the molecular structure of a newly discovered drug compound, for instance, can provide essential insights into its therapeutic potential and biological interactions, which is crucial for enhancing our understanding of related scientific literature.
To evaluate our model’s ability to understand molecular structures, we carried out a series of tasks involving molecules, polymers, and Markush structures, common in chemistry and medicine studies. These tasks were designed to assess the model’s capability in deciphering information from representations of molecular structures. The execution of these tasks and their respective results are detailed in Table 4.
The Uni-SMART model showed outstanding performance in molecular structure understanding tasks, particularly in processing complex molecular structures within the scientific literature. This is evident from the results presented in Table 4. For example, in the "Tag to Molecule" task, Uni-SMART achieved an average similarity score of 0.275, significantly outperforming other models which struggled with this task (GPT-4 and Gemini achieved a score of 0, while GPT-3.5 only reached 0.007). In the "Markush to Molecule" task, its average similarity score of 0.629 was also markedly higher than its competitors. These results clearly indicate Uni-SMART’s advantages in understanding molecular structures in scientific literature.
However, despite its overall excellent performance across most tasks, Uni-SMART was slightly inferior to GPT-3.5 in the "Affinity Data Extraction" task. This task involves extracting affinity tables from scientific literature, requiring not only an understanding of complex, domain-specific language but also a deep analysis to match text with molecular structures and other modalities. This suggests that while Uni-SMART has strong capabilities in molecular structure understanding, there is still room for improvement in handling tasks requiring detailed information retrieval and cross-modal matching, indicating potential areas for further optimization.
Table 4: Evaluation on molecule data.
Domain
Task
Metric
Uni-SMART
GPT-4
GPT-3.5
Gemini
Drug
Discovery
Affinity Data Extraction
Value Recall
0.320
0.274
0.359
0.175
Tag to Molecule
Mean Similarity
0.275
0.000
0.007
0.000
Molecule in Document
Accuracy
0.889
0.022
0.489
0.500
Markush to Molecule
Mean Similarity
0.629
0.545
0.459
0.199
Organic
Materials
OLED Property Extraction
Value Recall
0.526
0.520
0.479
0.394
3.4 Chemical Reactions
Understanding chemical reactions is crucial in the realm of scientific literature, particularly within chemistry and its related disciplines [16]. The ability to accurately parse and analyze the details of chemical reactions enhances readers’ grasp of complex material, enabling a deeper understanding of experimental results and theoretical discussions.
To assess Uni-SMART’s capability to comprehend chemical reactions, we designed two tasks aimed at challenging the model’s proficiency in interpreting chemical reactions. These tasks focus on understanding the reactants, products, and conditions of chemical reactions, as well as grasping the underlying mechanisms and clarifying the significance of these reactions within a broader scientific context. The results of these tasks, and a comparative analysis of Uni-SMART’s performance against other leading models, are summarized in Table 5.
Uni-SMART exhibited outstanding performance in "Reaction QA" within the drug discovery domain and "Reaction Mechanism QA" within the organic materials domain, achieving an accuracy of 0.4 and 0.445, respectively. These results significantly surpass those of GPT-4, GPT-3.5, and Gemini models. Such achievements highlight Uni-SMART’s considerable advantage in parsing chemical reactions, potentially attributed to its access to richer, more specialized training data sources, effective preprocessing of chemical expressions, and architecture designed specifically for multimodal data. Despite these strengths, there remains room for improvement in Uni-SMART’s comprehension of chemical reactions. Future enhancements may include strengthening the model’s understanding of reaction details and expanding the training dataset to encompass a broader range of reaction types and mechanisms. These efforts are expected to further augment Uni-SMART’s applicability in reading and interpreting scientific literature.
Table 5: Evaluation on chemical reactions data.
Domain
Task
Metric
Uni-SMART
GPT-4
GPT-3.5
Gemini
Drug Discovery
Reaction QA
Accuracy
0.400
0.200
0.000
0.133
Organic Materials
Reaction Mechanism QA
Accuracy
0.445
0.382
0.345
0.373
4 Application
In this section, we showcase how Uni-SMART significantly improves the efficiency and quality of scientific research. Through several selected case studies, we will present the specific applications of Uni-SMART in the scientific domain. These cases allow us to qualitatively assess the model’s capabilities in addressing various research challenges. For more details and updates, please refer to the Uni-SMART website at http://uni-smart.dp.tech/.
4.1 Patent infringement analysis
Figure 3: Patent infringement analysis.
In today’s research and industrial domains, the correct understanding and application of patent information have become increasingly crucial, especially in the fields of chemistry and drug development [17]. Accurately determining whether a compound is protected by existing patents is essential for avoiding potential patent infringements. However, this task presents numerous technical challenges, particularly the need for precise analysis of complex chemical structures and their comparison with descriptions in patent documents. This requires not only a deep understanding of chemical structures but also the ability to process and analyze information from different modalities, such as chemical structure and patent text descriptions.
In this case study, we utilize Uni-SMART to analyze whether a specific chemical structure is protected by a particular patent. The process involves receiving the patent document and compound uploaded by the user, analyzing the molecular skeleton and substituents of the compound, and then comparing this information with the range given in the patent document.
Figure 4: Understanding the Metal Injection Molding (MIM) Process.
As shown in Figure 3, Uni-SMART analyzes the provided compound structure and accurately determines whether the compound falls within the patent protection scope by considering multiple factors, such as the molecular skeleton and substituents. This process not only demonstrates Uni-SMART’s efficient handling of cross-modal information between chemical structure and patent text but also highlights its strong capability in understanding and applying patent information. Applying Uni-SMART in practical scenarios like patent infringement protection can help users avoid potential patent risks, which is of significant importance for advancing scientific research and industrial development.
4.2 Chart analysis
In scientific literature, charts serve as a crucial tool for conveying complex data and experimental results. They play an essential role in the deep understanding of technical details by presenting data trends in an intuitive form. However, the effective understanding and analysis of charts in scientific literature encounter significant challenges. These include but are not limited to, the simultaneous parsing of visual elements and textual descriptions within charts, the accurate identification of data points, understanding the scientific processes behind the data, and performing subsequent complex calculations and logical reasoning.
In this case study, Uni-SMART was utilized to analyze a chart showing the temperature control curve during the Metal Injection Molding (MIM) process. As illustrated in Figure 4, Uni-SMART accurately described the temperature control curve of the MIM process, precisely identifying the key data points in the chart. Following user instructions, it successfully converted this information into a tabular format. Furthermore, the model demonstrated its powerful capability in performing mathematical calculations and logical reasoning, especially in calculating the quenching rate.
5 Discussion and Conclusion
In this report, we introduce Uni-SMART for in-depth understanding of multimodal information within the scientific literature. Through rigorous quantitative evaluation, Uni-SMART demonstrates significant performance gain in interpreting and analyzing multimodal contents in scientific documents, such as tables, charts, molecular structures, and chemical reactions, compared with other leading competitors. The success of Uni-SMART lies in its innovative cyclic iterative process that continuously refines its multimodal understanding capabilities, leveraging a robust dataset and a combination of multimodal learning, supervised fine-tuning, user feedback, expert annotation, and data enhancement to achieve superior performance in scientific literature analysis.
Beyond quantitative assessment, we are particularly excited about Uni-SMART’s potential to address scientific challenges through practical applications. From patent infringement analysis to complex material science chart interpretation, Uni-SMART’s cross-modal understanding capabilities offer new perspectives and tools for research and technological development, showcasing its potential to facilitate research processes and accelerate discovery phases.
Despite Uni-SMART’s strong ability in multimodal scientific literature understanding, we acknowledge that there is still room for improvement. This includes enhancing the model’s understanding of highly complex and specialized content, as well as reducing hallucinations. We believe that through continuous research and development, these limitations will be addressed, making Uni-SMART an even more powerful and flexible tool for scientific research assistance.
In summary, the research and development of Uni-SMART mark a significant advancement in the field of multimodal scientific literature understanding. By providing scientists and researchers with an efficient tool for deep understanding and analysis of scientific documents, Uni-SMART not only facilitates the accumulation and innovation of scientific knowledge but also paves the way for future scientific work, technological development, and potential commercial applications. As we continue to improve and expand Uni-SMART, we look forward to its greater role in promoting scientific discovery and technological innovation.
Appendix
Appendix A Experiment Setup
For the evaluation of tables, charts, and chemical reactions, we conducted tests by uploading articles containing such content in PDF format and then posing questions to the models. Given the varying abilities of different models to process PDFs, we employed distinct methods to provide the PDFs.
For GPT-4, we utilized the web-based ChatGPT4 interface. We directly uploaded the original PDF files to the chat interface and posed queries, thereby leveraging OpenAI’s built-in PDF processing capabilities.
For GPT-3.5, we processed the PDFs using PyPDF2 to convert them into plain text, which was then inputted into the models.
For Gemini, given its proficiency in concurrently processing text and images, we initially employed PyPDF2 to extract text from the PDFs. Subsequently, we utilized PyMuPDF to retrieve images within the documents. Both text and images were then relayed to the model.
For Uni-SMART, owing to its capability to directly read PDF files, we simply uploaded the original PDF files and posed questions to the model.

### v2

@@
 However, existing LLMs have their own limits.
-Scientific literature often includes a wide range of multimodal elements, such as molecular structure, tables, and charts, which are hard for text-focused LLMs to understand and analyze.
+Scientific literature often includes a wide range of multimodal elements, such as tables, charts, and molecule, which are hard for text-focused LLMs to understand and analyze.
 This issue points to the urgent need for new solutions that can fully understand and analyze multimodal content in scientific literature.
 To answer this demand, we present Uni-SMART (Universal Science Multimodal Analysis and Research Transformer), an innovative model designed for in-depth understanding of multimodal scientific literature.
-Through rigorous quantitative evaluation across several domains, Uni-SMART demonstrates superior performance over leading text-focused LLMs.
+Through rigorous quantitative evaluation across several domains, Uni-SMART demonstrates superior performance over other text-focused LLMs.
 Furthermore, our exploration extends to practical applications, including patent infringement detection and nuanced analysis of charts.
@@
 Such ability not only augments automated and precise information extraction but also enriches the interaction between researchers and the vast expanse of scientific knowledge.
-To rigorously assess the multimodal capabilities of Uni-SMART, a comparative analysis was conducted against leading LLMs, such as Gemini, GPT-3.5, and GPT-4.
+To rigorously assess the multimodal capabilities of Uni-SMART, a comparative analysis was conducted against several LLMs, such as GPT-4o, Gemini, and Claude3.
 Our assessment targets several data types critical to the comprehension of scientific documents: tables, charts, molecular structures, and chemical reactions.
@@
 In the following sections, Section 2 details Uni-SMART’s data sources and iterative training approach.
-Subsequently, Section 3 – Evaluation, presents detailed comparisons of Uni-SMART with leading LLMs across a variety of modalities, showcasing its advanced capabilities in multimodal data interpretation.
+Subsequently, Section 3 – Evaluation, presents detailed comparisons of Uni-SMART with several LLMs across a variety of modalities, showcasing its advanced capabilities in multimodal data interpretation.
 Then, Section 4 – Application showcases some specific applications of Uni-SMART in the scientific domain.
@@
 In particular, It adopts a cyclical, iterative approach to enhance its multimodal understanding capabilities, comprising the following key components:
-  • Multimodal Learning: During the initialization phase, the model is trained with a limited set of multimodal data to recognize and extract diverse information elements from scientific literature, outputting these in a sequential format that incorporates both textual and multimodal data.
-  • LLM SFT: The output sequences, along with corresponding question-answer pairs, are utilized for supervised fine-tuning (SFT) of LLM, boosting its proficiency in handling multimodal content.
-  • User Feedback: The SFT-enhanced LLM is deployed in real applications, facilitating the collection of user feedback.
+  • Multimodal Learning: During the initialization phase, the parsing model is trained with a limited set of multimodal data to recognize and extract diverse information elements from scientific literature.
+The output is formatted in a custom text format, similar to html format, to effectively represent multimodal elements.
+  • LLM SFT: A series of valuable queries are constructed, particularly in multimodal scenarios.
+With Multimodal Retrieval-Augmented Generation (RAG), relevant content is recalled from the literature based on queries.
+Answers are built based on the queries and retrieved content.
+The query-answer data is then used to fine-tune the LLM.
+This process helps the LLM adapt to the custom input format and improves its ability to follow instructions.
+  • User Feedback: The parsing model and SFT-enhanced LLM are deployed in real applications, facilitating the collection of user feedback.
 Samples receiving positive feedback are subsequently filtered and incorporated into the data enhancement, while those with negative feedback are subject to expert annotation before being integrated into the data enhancement process.
-  • Expert Annotation: Samples with negative feedback are meticulously annotated by human experts.
-This step ensures that the model learns from its mistakes, with semi-automated tools assisting in this process to enhance efficiency.
-Negative feedback is examined to identify challenges in multimodal recognition or the language model’s reasoning abilities.
-This analysis helps determine which multimodal elements and types of question are more challenging for the model, facilitating focused improvements.
+  • Expert Annotation: Samples with negative feedback are carefully annotated by human experts.
+This step ensures that these models learns from their mistakes, with semi-automated tools assisting in this process to enhance efficiency.
+Negative feedback cases typically fall into the following categories:
+    1. 1.
+      Multimodal element recognition errors;
+    2. 2.
+      Recall content errors;
+    3. 3.
+      Poor instruction-following.
+    Detailed analysis of these error types facilitates targeted improvements.
   • Data Enhancement: Finally, the annotated data, along with partial samples that received positive feedback, are added to the training dataset for data enhancement.
-This iterative process progressively improves Uni-SMART’s performance, enriching its training dataset and enhancing its multimodal understanding capabilities.
-Such cyclical pipeline significantly enhances Uni-SMART’s performance in a variety of challenging tasks, such as information extraction, complex element identification, scientific literature understanding/analysis, and multimodal understanding/reasoning.
+The pipeline is optimized based on different types of negative feedback:
+    1. 1.
+      Multimodal element recognition errors: expand the training data for the parsing model;
+    2. 2.
+      Recall content errors: optimize the Multimodal RAG scheme;
+    3. 3.
+      Poor instruction-following: enlarge and refine the dataset used for LLM SFT.
+This iterative process is repeated to continually optimize Uni-SMART’s overall performance.
+And significantly enhances Uni-SMART’s performance in a variety of challenging tasks, such as information extraction, complex element identification, scientific literature understanding/analysis, and multimodal understanding/reasoning.
 Figure 2: Uni-SMART model architecture.
 3 Evaluation
-In this section, we perform a detailed quantitative evaluation on the capabilities of Uni-SMART and leading LLMs across various modalities.
+In this section, we perform a detailed quantitative evaluation on the capabilities of Uni-SMART and various available LLMs across modalities.
 Table 1 presents the statistics of the benchmarks.
 For details of the specific evaluation tasks, please refer to SciAssess [10].
-And the experimental setup for each model is detailed in Appendix A.
-Table 1: Dataset statistics.
-Modality
-Domain
-Task
-Question #
-Table
-Alloy Materials
-Composition Extraction
-55
-Drug Discovery
-Affinity Data Extraction
-15
-Organic Materials
-Electrolyte Solubility data extraction
-8
-Electrolyte Table QA
-48
-Polymer Property Extraction
-15
-OLED Property Extraction
-13
-Chart
-Alloy Materials
-Alloy ChartQA
-15
-Biology
-Biology ChartQA
-15
-Drug Discovery
-Drug ChartQA
-15
-Organic Materials
-Polymer ChartQA
-15
-Molecule
-Drug Discovery
-Affinity Data Extraction
-15
-Tag to Molecule
-41
-Markush to Molecule
-9
-Organic Materials
-OLED Property Extraction
-13
-Chemical Reaction
-Drug Discovery
-Reaction QA
-15
-Organic Materials
-Reaction Mechanism QA
-22
+For our experiments, different methods are employed depending on the source of the models.
+Closed-source models are accessed via API calls, while open-source models are obtained from Hugging Face, deployed, and tested.
+Evaluation tasks that require an article as context involve converting the PDF content to text for input into the LLMs.
+If an LLM offers a built-in PDF parsing interface, we utilize this feature; otherwise, PyPDF2 is used for PDF parsing.
+Notably, Uni-SMART can directly read PDF files, allowing us to upload the original documents and pose questions directly to the model.
+Due to input length limitations of the LLMs, tasks that require the full text of an article as context are executed in a zero-shot manner, whereas tasks that do not require such extensive context are tested using few-shot approaches.
+The final results of these experiments are systematically presented in Table 2, illustrating the performance of each model across the evaluated tasks.
+Table 1: Statistics of evaluation datasets
+Modality Domain Task Question Type Metric # Question
+Table Alloy Materials Composition Extraction Table Extraction Table Recall 244
+Drug Discovery Affinity Extraction Table Extraction Table Recall 40
+Organic Materials Electrolyte Table QA Multi Choices Accuracy 100
+OLED Property Extraction Table Extraction Table Recall 13
+Polymer Property Extraction Table Extraction Table Recall 109
+Solubility Extraction Table Extraction Table Recall 100
+Chart Alloy Materials Alloy Chart QA Multi Choices Accuracy 15
+Biology Biology Chart QA Multi Choices Accuracy 99
+Drug Discovery Drug Chart QA Multi Choices Accuracy 15
+Organic Materials Polymer Chart QA Multi Choices Accuracy 15
+Molecule Drug Discovery Affinity Data Extraction Table Extraction Table Recall 40
+Tag to Molecule Molecule Generation Molecule Similarity 50
+Markush to Molecule Molecule Generation Molecule Similarity 37
+Molecule in Document YES/NO Accuracy 50
+Organic Materials OLED Property Extraction Table Extraction Table Recall 13
+Reaction Drug Discovery Reaction QA Multi Choices Accuracy 95
+Organic Materials Reaction Mechanism QA Multi Choices Accuracy 22
+Table 2: Performance Comparison of LLMs on Diverse Tasks.
+Modality Task Uni-SMART GPT-4o GPT-4 GPT-3.5 Moonshot Claude3 Doubao Gemini Llama3 DeepSeek Qwen2 Command $\text{R}^{{\text{+}}}$
+Table Composition Extraction 0.511 0.484 0.458 0.112 0.127 0.495 0.304 0.239 0.212 0.389 0.423 0.128
+Affinity Extraction 0.200 0.072 0.042 0.025 0.040 0.097 0.050 0.040 0.064 0.017 0.075 0.043
+Electrolyte Table QA 0.850 0.890 0.790 0.370 0.670 0.870 0.710 0.880 0.460 0.720 0.620 0.450
+OLED Property Extraction 0.490 0.336 0.406 0.201 0.037 0.477 0.259 0.093 0.263 0.292 0.392 0.234
+Polymer Property Extraction 0.705 0.692 0.681 0.329 0.705 0.629 0.514 0.606 0.536 0.652 0.636 0.171
+Solubility Extraction 0.468 0.435 0.440 0.410 0.363 0.426 0.371 0.397 0.399 0.432 0.400 0.351
+Chart Alloy Chart QA 0.933 0.533 0.600 0.333 0.333 0.400 0.467 0.667 0.467 0.333 0.400 0.200
+Biology Chart QA 0.616 0.580 0.480 0.390 0.545 0.505 0.480 0.616 0.520 0.545 0.515 0.535
+Drug Chart QA 0.600 0.333 0.400 0.067 0.400 0.200 0.533 0.533 0.400 0.400 0.400 0.533
+Polymer Chart QA 0.933 0.800 0.667 0.400 0.800 0.467 0.867 0.800 0.867 0.733 0.933 0.800
+Molecule Affinity Extraction 0.200 0.072 0.042 0.025 0.040 0.097 0.050 0.040 0.064 0.017 0.075 0.043
+Tag to Molecule 0.188 0.040 0.022 0.000 0.016 0.035 0.094 0.169 0.034 0.014 0.000 0.031
+Markush to Molecule 0.686 0.642 0.654 0.431 0.504 0.675 0.239 0.526 0.491 0.470 0.379 0.376
+Molecule in Document 0.720 0.580 0.700 0.500 0.460 0.480 0.560 0.640 0.680 0.460 0.460 0.460
+OLED Property Extraction 0.490 0.336 0.406 0.201 0.037 0.477 0.259 0.093 0.263 0.292 0.392 0.234
+Reaction Reaction QA 0.768 0.705 0.674 0.442 0.253 0.663 0.442 0.305 0.611 0.368 0.442 0.316
+Reaction Mechanism QA 0.682 0.545 0.636 0.455 0.545 0.455 0.636 0.727 0.500 0.545 0.591 0.591
 3.1 Table
@@
 The utilization of table data spans a wide array of applications, such as trend analysis, which can reveal developmental trajectories in research fields, and comparative studies, which can elucidate differences in experimental outcomes under varying research methodologies or conditions [12].
-To assess the table understanding capabilities of our model, Uni-SMART, compared to other leading models, we designed a diverse set of tasks across different domains.
-These tasks were specifically tailored to evaluate how well each model could interpret and extract information from tables, with the results presented in Table 2.
+To assess the table understanding capabilities of our model, Uni-SMART, compared to other LLMs, we designed a diverse set of tasks across different domains.
+These tasks were specifically tailored to evaluate how well each model could interpret and extract information from tables.
 In our thorough assessment, the Uni-SMART showcased its exceptional proficiency in understanding and extracting table data from the scientific literature.
-Among six assessed tasks, it surpassed other models in four and delivered a competitive performance in the remainder.
-Notably, in the "Electrolyte Table QA" and "Polymer Property Extraction" tasks, Uni-SMART achieved Value Recall scores of 0.674 and 0.869, respectively, significantly outperforming its counterparts.
-These results highlight Uni-SMART’s exceptional skill in processing table data.
-However, Uni-SMART’s performance in the "Composition Extraction" task for Alloy Materials showed no significant difference compared to other models, only slightly leading.
-To further enhance its effectiveness, a possible approach could be to strengthen the model’s ability to learn domain-specific knowledge.
-Moreover, all models under-performed in the "Affinity Data Extraction" task within Drug Discovery and the "Electrolyte Solubility data extraction" task in Organic Materials, indicating the high complexity of these tasks.
-Such challenges might be attributed to the diversity of table formats and the presence of complex nested structures, such as tables embedding molecular structures.
-To navigate these difficulties, future directions could involve the development of more sophisticated table data comprehension technologies and the refinement of data preprocessing methods.
-Table 2: Evaluation on table data.
-Domain
-Task
-Metric
-Uni-SMART
-GPT-4
-GPT-3.5
-Gemini
-Alloy Materials
-Composition Extraction
-Value Recall
-0.451
-0.420
-0.437
-0.443
-Drug Discovery
-Affinity Data Extraction
-Value Recall
-0.320
-0.274
-0.359
-0.175
-Organic
-Materials
-Electrolyte Solubility data extraction
-Value Recall
-0.282
-0.365
-0.393
-0.390
-Electrolyte Table  QA
-Value Recall
-0.674
-0.163
-0.140
-0.233
-Polymer Property Extraction
-Value Recall
-0.869
-0.762
-0.286
-0.506
-OLED Property Extraction
-Value Recall
-0.526
-0.520
-0.479
-0.394
+Among six assessed tasks, it surpassed other models in five and delivered a competitive performance in the remainder.
+Notably, in the "Composition Extraction" and "Solubility Extraction" tasks, Uni-SMART achieved "Table Recall" of 0.511 and 0.468, respectively, significantly outperforming its counterparts.
+The objectives of these two tasks are to extract necessary information from tables within articles and organize it into a specified format.
+These results indicate that Uni-SMART excels in handling and understanding tabular data, particularly in information extraction and formatting tasks.
+However, in the Electrolyte Table QA task, Uni-SMART’s performance was slightly below the state-of-the-art model (GPT-4o).
+In this task, LLMs are required to answer questions about details in the table with the question type of multiple-choice, indicating that Uni-SMART still has room for improvement in understanding the details within tables.
+Moreover, all models under-performed in the "Affinity Extraction" task, indicating the high complexity of these tasks.
+This task requires LLMs to organize molecules and corresponding experimental result from articles.
+It necessitates the model’s ability to integrate molecules with table information and demands strong long-context understanding capabilities.
+To better address these issues in the future, a possible direction is to enhance LLMs’ abilities in long context comprehension and information matching.
 3.2 Chart
@@
 These tasks rigorously evaluated models’ ability to analyze and clarify data presented in charts, with a particular focus on identifying trends and extracting meaningful insights.
-Analysis of the results presented in Table 3 indicates that Uni-SMART significantly outperforms existing models like GPT-4, GPT-3.5, and Gemini in ChartQA tasks across diverse scientific domains, particularly excelling in the areas of alloy materials and organic materials, with leading scores of 0.667 and 0.733, respectively.
-However, despite Uni-SMART’s leading position, its advantage narrows in the domains of drug discovery and biology, where all models struggle to achieve high performance.
-This limitation is likely due to the complexity and specificity of charts in these fields, which demand higher levels of understanding and adaptability from the models.
-To overcome these challenges, future work should focus on augmenting the training dataset with a wider and more sophisticated assortment of chart data, aiming to enhance the model’s analytical depth and accuracy.
-Furthermore, exploring refined model architectures and specialized training methodologies, specifically designed to meet the complex requirements of chart analysis, is crucial for significantly elevating Uni-SMART’s performance in such demanding tasks.
-Table 3: Evaluation on chart data.
-Domain
-Task
-Metric
-Uni-SMART
-GPT-4
-GPT-3.5
-Gemini
-Alloy Materials
-Alloy ChartQA
-Accuracy
-0.667
-0.400
-0.267
-0.333
-Biology
-Biology ChartQA
-Accuracy
-0.467
-0.333
-0.400
-0.400
-Drug Discovery
-Drug ChartQA
-Accuracy
-0.467
-0.400
-0.333
-0.400
-Organic Materials
-Polymer ChartQA
-Accuracy
-0.733
-0.667
-0.333
-0.007
-3.3 Molecular Structure
-In scientific literature, molecular structures hold fundamental importance, especially within fields such as chemistry, pharmacology, and biology.
-The ability to comprehend molecular structures is crucial for analyzing research outcomes, predicting chemical behaviours, and innovating novel compounds [14, 15].
+Analysis of the results presented in Figure 3 indicates that Uni-SMART significantly outperforms existing models in Chart QA tasks across diverse scientific domains, particularly excelling in the tasks of Alloy Chart QA and Drug Chart QA, with leading scores of 0.933 and 0.600, respectively.
+However, both in the Biology Chart QA, where the Gemini model scored 0.616, and in the Polymer Chart QA task, where Qwen2 also achieved a top score of 0.933—matching Uni-SMART’s performance in both cases—it is evident that these models have strong capabilities in analyzing scientific charts.
+These results underscore the intense competition among different models in the context of chart understanding.
+The future direction for development likely focuses on enhancing accuracy and applicability across various scientific domains.
+3.3 Molecule
+In scientific literature, molecule or molecular structures hold fundamental importance, especially within fields such as chemistry, pharmacology, and biology.
+Comprehending molecular structures is crucial for analyzing research outcomes, predicting chemical behaviors, and innovating novel compounds [14, 15].
 Deciphering the molecular structure of a newly discovered drug compound, for instance, can provide essential insights into its therapeutic potential and biological interactions, which is crucial for enhancing our understanding of related scientific literature.
-To evaluate our model’s ability to understand molecular structures, we carried out a series of tasks involving molecules, polymers, and Markush structures, common in chemistry and medicine studies.
+To evaluate our model’s ability to understand molecular structures, we carried out a series of tasks involving molecules, polymers, and Markush structures, which are common in chemistry and medicine studies.
 These tasks were designed to assess the model’s capability in deciphering information from representations of molecular structures.
-The execution of these tasks and their respective results are detailed in Table 4.
 The Uni-SMART model showed outstanding performance in molecular structure understanding tasks, particularly in processing complex molecular structures within the scientific literature.
-This is evident from the results presented in Table 4.
-For example, in the "Tag to Molecule" task, Uni-SMART achieved an average similarity score of 0.275, significantly outperforming other models which struggled with this task (GPT-4 and Gemini achieved a score of 0, while GPT-3.5 only reached 0.007).
-In the "Markush to Molecule" task, its average similarity score of 0.629 was also markedly higher than its competitors.
-These results clearly indicate Uni-SMART’s advantages in understanding molecular structures in scientific literature.
-However, despite its overall excellent performance across most tasks, Uni-SMART was slightly inferior to GPT-3.5 in the "Affinity Data Extraction" task.
-This task involves extracting affinity tables from scientific literature, requiring not only an understanding of complex, domain-specific language but also a deep analysis to match text with molecular structures and other modalities.
-This suggests that while Uni-SMART has strong capabilities in molecular structure understanding, there is still room for improvement in handling tasks requiring detailed information retrieval and cross-modal matching, indicating potential areas for further optimization.
-Table 4: Evaluation on molecule data.
-Domain
-Task
-Metric
-Uni-SMART
-GPT-4
-GPT-3.5
-Gemini
-Drug
-Discovery
-Affinity Data Extraction
-Value Recall
-0.320
-0.274
-0.359
-0.175
-Tag to Molecule
-Mean Similarity
-0.275
-0.000
-0.007
-0.000
-Molecule in Document
-Accuracy
-0.889
-0.022
-0.489
-0.500
-Markush to Molecule
-Mean Similarity
-0.629
-0.545
-0.459
-0.199
-Organic
-Materials
-OLED Property Extraction
-Value Recall
-0.526
-0.520
-0.479
-0.394
+For example, in the "Molecule in Document" task, Uni-SMART achieved an accuracy of 0.720, significantly outperforming other models.
+This task requires LLMs to determine whether a molecular structure has appeared in the literature, demonstrating Uni-SMART’s strong capabilities in accurate molecule identification and information retrieval.
+Such achievements highlight Uni-SMART’s considerable advantage in parsing molecular structure, potentially attributed to its access to richer, more specialized training data sources, effective preprocessing of molecular structure, and architecture designed specifically for multimodal data.
+Despite these strengths, there remains room for improvement in Uni-SMART’s comprehension of molecular structure.
+Future enhancements may include strengthening the model’s understanding of molecule details and expanding the training dataset to encompass a broader range of molecule types.
+These efforts are expected to further augment Uni-SMART’s applicability in reading and interpreting scientific literature, especially those involving molecular structure.
 3.4 Chemical Reactions
-Understanding chemical reactions is crucial in the realm of scientific literature, particularly within chemistry and its related disciplines [16].
+Understanding chemical reactions is crucial in scientific literature, particularly within chemistry and its related disciplines [16].
 The ability to accurately parse and analyze the details of chemical reactions enhances readers’ grasp of complex material, enabling a deeper understanding of experimental results and theoretical discussions.
@@
 These tasks focus on understanding the reactants, products, and conditions of chemical reactions, as well as grasping the underlying mechanisms and clarifying the significance of these reactions within a broader scientific context.
-The results of these tasks, and a comparative analysis of Uni-SMART’s performance against other leading models, are summarized in Table 5.
-Uni-SMART exhibited outstanding performance in "Reaction QA" within the drug discovery domain and "Reaction Mechanism QA" within the organic materials domain, achieving an accuracy of 0.4 and 0.445, respectively.
-These results significantly surpass those of GPT-4, GPT-3.5, and Gemini models.
-Such achievements highlight Uni-SMART’s considerable advantage in parsing chemical reactions, potentially attributed to its access to richer, more specialized training data sources, effective preprocessing of chemical expressions, and architecture designed specifically for multimodal data.
-Despite these strengths, there remains room for improvement in Uni-SMART’s comprehension of chemical reactions.
-Future enhancements may include strengthening the model’s understanding of reaction details and expanding the training dataset to encompass a broader range of reaction types and mechanisms.
-These efforts are expected to further augment Uni-SMART’s applicability in reading and interpreting scientific literature.
-Table 5: Evaluation on chemical reactions data.
-Domain
-Task
-Metric
-Uni-SMART
-GPT-4
-GPT-3.5
-Gemini
-Drug Discovery
-Reaction QA
-Accuracy
-0.400
-0.200
-0.000
-0.133
-Organic Materials
-Reaction Mechanism QA
-Accuracy
-0.445
-0.382
-0.345
-0.373
+Uni-SMART exhibited outstanding performance in "Reaction QA" task, achieving an accuracy of 0.768.
+However, in the "Reaction Mechanism QA" task, Uni-SMART’s performance was slightly below that of the Gemini.
+This task requires LLMs to accurately parse and interpret the step-by-step mechanisms of chemical reactions from the article, indicating that Uni-SMART still has room for improvement in understanding complex reactions.
+Figure 3: Comparative Performance of LLMs Across Different Modalities.
 4 Application
@@
 In this report, we introduce Uni-SMART for in-depth understanding of multimodal information within the scientific literature.
-Through rigorous quantitative evaluation, Uni-SMART demonstrates significant performance gain in interpreting and analyzing multimodal contents in scientific documents, such as tables, charts, molecular structures, and chemical reactions, compared with other leading competitors.
+Through rigorous quantitative evaluation, Uni-SMART demonstrates significant performance gain in interpreting and analyzing multimodal contents in scientific documents, such as tables, charts, molecular structures, and chemical reactions, compared with other competitors.
 The success of Uni-SMART lies in its innovative cyclic iterative process that continuously refines its multimodal understanding capabilities, leveraging a robust dataset and a combination of multimodal learning, supervised fine-tuning, user feedback, expert annotation, and data enhancement to achieve superior performance in scientific literature analysis.
@@
 As we continue to improve and expand Uni-SMART, we look forward to its greater role in promoting scientific discovery and technological innovation.
-Appendix
-Appendix A Experiment Setup
-For the evaluation of tables, charts, and chemical reactions, we conducted tests by uploading articles containing such content in PDF format and then posing questions to the models.
-Given the varying abilities of different models to process PDFs, we employed distinct methods to provide the PDFs.
-For GPT-4, we utilized the web-based ChatGPT4 interface.
-We directly uploaded the original PDF files to the chat interface and posed queries, thereby leveraging OpenAI’s built-in PDF processing capabilities.
-For GPT-3.5, we processed the PDFs using PyPDF2 to convert them into plain text, which was then inputted into the models.
-For Gemini, given its proficiency in concurrently processing text and images, we initially employed PyPDF2 to extract text from the PDFs.
-Subsequently, we utilized PyMuPDF to retrieve images within the documents.
-Both text and images were then relayed to the model.
-For Uni-SMART, owing to its capability to directly read PDF files, we simply uploaded the original PDF files and posed questions to the model.